# Revealing the nature of ultra-long period objects with space-based gravitational-wave interferometers

### v1

@@ 3.1 Signal-to-noise ratios
for speed of light $c$ and Newton’s constant $G$. We calculate expression (4) for sources listed in Tab. 1. The results are shown in Table 2 for two scenarios: (i) using the mean value of $\mathcal{M}$ from expression (1) along with best-fit distances, and (ii) an optimistic case where the masses and distances take their upper and lower confidence bounds, respectively. The latter scenario is indicated by bracketed values in the second ($T=1$ mth) and third ($T=4$ yr) columns of Tab. 2.
As expected visually from Fig. 1, taking $T\lesssim 4$ yr all but guarantees that CHIME J0630, GX J1627, and GPM J1839 would be identifiable in the LISA data stream. The former two should be visible within a month of first light. While hopes for other ULPs are low, there is some call for optimism for DA J1832 and A J1935 if the components are heavier than average, allowing them to reach significant SNRs when accounting for their sky locations. Because DA J1832 has shown contemporaneous X-ray bursting with radio pulsing (Wang et al., 2024), a magnetar progenitor is favoured over an WDMD (Suvorov, Dehman & Pons, 2025); a detection would thus insist on revisions to theoretical models.
@@
Table 2: SNRs for LISA for the systems listed in Tab. 1 for $T=1$ mth (middle column) or $T=4$ yr (right) taking either mean values of the chirp mass, expression 1, and distance or favourable extrema thereof (bracketed values).
Source name SNR (1 mth; max) SNR (4 yr; max)
CHIME J0630+25 667 ($>10^{3}$) $>10^{3}$ ($>10^{4}$)
GLEAM-X J1627 15.9 (48.6) 110 (337)
GPM J1839–10 2.22 (8.53) 15.4 (59.1)
DA J1832 0.42 (0.95) 2.92 (6.59)
ASKAP J1935 0.21 (0.45) 1.47 (3.08)
GCRT J1745–3009 0.04 (0.08) 0.30 (0.57)
ILT J1101+5521 0.14 (0.36) 0.98 (2.49)
GLEAM-X J0704 0.10 (0.19) 0.67 (1.28)
J1912–4410 0.02 (0.05) 0.13 (0.32)
ASKAP J1839 $<10^{-3}$ $(<10^{-3})$ 0.002 (0.006)
4 Conclusions and discussion
@@
ULPs remain mysterious, with some studies favouring isolated neutron stars and others favouring WDMDs (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). We have argued here that if ULPs are WDMDs, at least a handful of them should be visible to LISA and its sister spacecraft (Tab. 2). If electromagnetic data remain inconclusive, GW observations could help unveil their nature. Importantly, among the known ULPs, those brightest in the LISA band correspond to objects with uncertain classification (Fig. 1). It is worth commenting that neutron-star-M-dwarf binaries are not observationally excluded (Hurley-Walker et al., 2024) and could also represent (a subclass of) ULPs. For primaries with $M_{\star}\approx 1.4M_{\odot}$, $\mathcal{M}$ would be larger by $\approx 50\%$ compared to white dwarfs, boosting SNRs by a factor $\sim 2$.
@@
While some implications of (non-)detections were highlighted in Sec. 3, a wealth of physical processes may operate in ULPs. For instance, linking radio luminosities to electromotive losses implies a minimum strength for the primary’s magnetic field. Assuming a maser-like emission mechanism, some systems appear to require GigaGauss strengths (Qu & Zhang, 2025). While dynamo action may amplify an internal field (Schreiber et al., 2021), the magnetospheric field may not reflect the interior strength because magnetic burial – a process where field lines are equatorially advected by infalling plasma – will reduce the global dipole moment during accretion while creating strong multipoles (Suvorov & Melatos, 2020). Typical accretion rates in tight-binary polars are $\sim 10^{-10}M_{\odot}\text{ yr}^{-1}$ (Pala et al., 2020) implying that, for a C/O dwarf with $M_{\star}\sim 0.8M_{\odot}$, the accretion timescale matches that of Ohmic diffusion ($\tau_{\Omega}$) at depths corresponding to densities of $\sim 10^{5}\text{ g cm}^{-3}$ where Cumming (2002) estimates $\tau_{\Omega}\lesssim 1$ Gyr. This suggests a timeline for radio activation post detachment if the interaction zone lies at a polar latitude, with the reverse applying near the equator if magnetic flux is compressed there. Resistive relaxation simulations would help elucidate such interactions and their connections with diamagnetic screening, crystallisation, and convection.

### v2

@@ 3.1 Signal-to-noise ratios
for speed of light $c$ and Newton’s constant $G$. We calculate expression (9) for sources listed in Tab. 1. The results are shown in Table 2 for two scenarios: (i) using the maximum DM-inferred distances and lowest values of $\mathcal{M}$ considered in Sec. 2, and (ii) an optimistic case where the masses and distances take their upper and lower confidence bounds, respectively. The latter scenario is indicated by bracketed values in the second ($T=0.5$ yr) and third ($T=4$ yr) columns of Tab. 2.
As expected visually from Fig. 1, taking $T\lesssim 0.5$ yr practically guarantees that CHIME J0630 and GX J1627 would be identifiable in the LISA data stream if $P=P_{\rm orb}$. While hopes for other ULPs are low, there is some call for optimism for DA J1832 and A J1935 if the systems contain heavier stars, allowing them to reach significant SNRs when accounting for their sky locations. For instance, if A J1935 contains a neutron star and a reasonably heavy companion, its well-constrained sky location could allow for a boost to the SNR by a factor $\sim 2$, which would give a final value of $\text{SNR}(4\text{yr})\sim 12$ which would be sufficient to claim a confident detection. Combining data from LISA with Taiji and TianQin could also lead to non-negligible boosts in SNR for this and other sources, perhaps enabling detections in marginal scenarios like a WDWD for DA J1832.
@@
Table 2: SNRs for LISA for the systems listed in Tab. 1 for $T=0.5$ yr (middle column) or $T=4$ yr (right) in either the least (largest distances and smallest chirp masses) or most (bracketed values) optimistic scenarios.
Source name SNR (0.5 yr; max) SNR (4 yr; max)
CHIME J0630+25 84.0 (233) 238 (660)
GLEAM-X J1627 31.0 (221.7) 87.8 (627)
GPM J1839–10 0.27 (0.58) 0.77 (1.65)
DA J1832 0.42 (4.33) 1.18 (12.3)
ASKAP J1935 0.22 (2.03) 0.63 (5.73)
GCRT J1745–3009 0.05 (0.50) 0.14 (1.41)
ILT J1101+5521 0.13 (0.22) 0.37 (0.62)
GLEAM-X J0704 0.05 (0.09) 0.15 (0.25)
ASKAP J1839 $<10^{-3}$ (0.004) $<10^{-3}$ (0.01)
4 Conclusions
@@
ULPs remain mysterious, with some studies favouring isolated neutron stars and others favouring WDMDs or compact binaries (see, e.g., Rea et al., 2022; Coti Zelati & Borghese, 2024). We have argued here that if the pulsation period in ULPs matches that of the orbit or close to it, at least a few should be visible to LISA and its sister spacecraft (Tab. 2). If electromagnetic data remain inconclusive, GW observations could help unveil their nature. Importantly, among the known ULPs, those that may be brightest in the LISA band correspond to objects with uncertain classification (Fig. 1).
@@
While some implications of (non-)detections were highlighted in Secs. 2 and 3, a wealth of physical processes may operate in ULPs. For instance, linking radio luminosities to electromotive losses implies a minimum strength for the primary’s magnetic field. Assuming a maser-like emission mechanism, some systems appear to require GigaGauss strengths (Qu & Zhang, 2025). While dynamo action may amplify an internal field (Schreiber et al., 2021), the magnetospheric field may not reflect the interior strength because magnetic burial – a process where field lines are equatorially advected by infalling plasma – will reduce the global dipole moment during accretion while creating strong multipoles (Suvorov & Melatos, 2020). Typical accretion rates in tight-binary polars are $\sim 10^{-10}M_{\odot}\text{ yr}^{-1}$ (Pala et al., 2020) implying that, for a C/O dwarf with $M_{\star}\sim 0.8M_{\odot}$, the accretion timescale matches that of Ohmic diffusion ($\tau_{\Omega}$) at depths corresponding to densities of $\sim 10^{5}\text{ g cm}^{-3}$ where Cumming (2002) estimates $\tau_{\Omega}\lesssim 1$ Gyr. This suggests a timeline for radio activation post detachment if the interaction zone lies at a polar latitude, with the reverse applying near the equator if magnetic flux is compressed there. Resistive relaxation simulations would help elucidate such interactions and their connections with diamagnetic screening, crystallisation, and convection.